A student stated that the solution for the inequality 6x - 2 – 17x > -35 is x > -33 Is the student correct? Why or why not? Show how you would solve this inequality.
Answer by lwsshak3(11628)

A student stated that the solution for the inequality 6x - 2 – 17x > -35 is x > -33 Is the student correct? Why or why not? Show how you would solve this inequality.
..
6x - 2 – 17x > -35
-11x-2>-35
-11x>-33
Divide both sides by -11 and switch inequality sign
x<3 (ans)
..
Check:
let x=3
6x - 2 – 17x=18-2-51=-35
To make the left side > right side, the left side must be less negative, that is, x<3